what is the intersection points of y= x^3-3x^2+x and y=x^2-3x
my concern is inputing x^3 into the quadratic equation

Answer by scott8148(6628)

the expressions are equal at the intersection(s)

x^3 - 3x^2 + x = x^2 - 3x

x^3 - 4x^2 + 4x = 0

factoring ___ x(x^2 - 4x + 4) = 0

still factoring ___ x(x - 2)^2 = 0

x = 0 and x = 2 (double root)

the intersection points are (0,0) and (2,-2)